b) Following are demand and price data for 10 randomly selected samples. Here, x denotes demand, in units, and y denotes price, in hundreds of dollars. Predict the prices for X=22 and 23. (2 marks)
X 16 18 19 20 21 25 24 25 24 21
Y 290 280 295 425 384 315 355 328 325 425
1
Expert's answer
2021-05-07T09:35:20-0400

1. Let us first determine "\sum x_i , \sum x^2_i , \sum y_i , \sum y^2_i , \sum x_iy_i"


"\sum x_i = 213 \\

\sum x^2_i = 4625 \\

\sum y_i = 3422 \\

\sum y^2_i = 1196690 \\

\sum x_iy_i = 73169"

n represents the sample size and thus n is equal to the number of ordered pairs.

n = 10

We can then determine the covariance using the formula

"s_{xy} = \frac{\sum x_iy_i - \frac{(\sum x_i)(\sum y_i)}{n}}{n -1} \\

= \frac{ 73169 - \frac{(213)(3422)}{10}}{10 -1} \\

= 31.15"

Let us next determine the sample variance s2 using the formula:

"s^2 = \frac{\sum x^2_i - \frac{(\sum x_i)^2}{n}}{n-1} \\

s^2_x = \frac{ 4625 - \frac{(213)^2}{10}}{10-1} = 9.78 \\

s^2_y = \frac{ 1196690 - \frac{(3422)^2}{10}}{10-1} = 2568.16"

The sample standard deviation is the square root of the population sample:

"s_x = \sqrt{s^2_x} = \sqrt{9.78} = 3.12 \\

s_y = \sqrt{s^2_y} = \sqrt{2568.16} = 50.67"

We can then determine the correlation coefficient r using the formula:

"r = \frac{s_{xy}}{s_x s_y} \\

r = \frac{31.15}{3.12 \times 50.67} = 0.197"

2. The relationship between the variables is weak, positive, linear relationship.

3. Next, we can determine the slope b using the formula:

"b = r\frac{s_y}{s_x} \\

b = 0.197 \times \frac{50.67}{3.12} = 3.199"

Next, we can determine the y-intercept a using the formula "a = \bar{y} -b \bar{x}" , where the sample mean is the sum of all values divided by the number of values.

"a = \bar{y} -b \bar{x} = \frac{\sum y_i}{n} -b \frac{\sum x_i}{n} \\

a = \frac{3422}{10} -3.199 \frac{213}{10} \\

= 342.2 -68.138 = 274.062"

Finally, we then obtain the regression line:

"y = a + bx = 274.062 + 3.199x \\

y(22) = 274.062 + 3.199 \times 22 = 344.44 \\

y(23) = 274.062 + 3.199 \times 23 = 347.64"